OpenStudy (anonymous):

Solve for x in the proportion . 3 6 ---- = ---- x+2 x-1 x = −5 x = 1 x = 4 x = 5

OpenStudy (zehanz):

You could do two things here: 1. Just plug the answers in to see which is right. 2. write the equation as\[3(x-1)=6(x+2)\]Then work out the brackets, collect terms (terms with x to the left side, the rest on the right) and solve for x.

OpenStudy (anonymous):

C??

OpenStudy (zehanz):

Well, the steps above lead to:\[3x-3=6x+12\]\[-3x=15\]\[x=\frac{ 15 }{ -3 }=-5\]so the right answer would be A.